### v1

@@ 5 Can Machine Unlearning Remove Poisons?
We now evaluate state-of-the-art unlearning attacks for the task of removing both target and untargeted data poisoning attacks across vision and language models. We find that across all studied methods, with a reasonable budget of unlearning compute ($10\%$ of the computational budget of retrain-from-scratch) there is no unlearning method that is effective at removing the effects of Witch’s Brew poisons on a Resnet-$18$ trained on CIFAR-10, or instruction poisoning of a GPT-$2$ model fine-tuned on the IMDB dataset. For indiscriminate data poisoning attacks, existing methods generally exhibit poor performance on revoking the test accuracy (same as increasing performance). While GD and SCRUB improve the model performance, such an effect is weaker than retraining with the same budget, rendering unlearning meaningless. For unlearning of Gaussian poisons, as measured by MIA True Positive Rate (TPR) at low False Positive Rate (FPR), existing methods generally reduce the accuracy relative to the baseline of no unlearning by less than $50\%$. For methods that do appear to unlearn reasonably well, NGD on Resnet-18 and SCRUB on GPT-2, there is a significant cost to accuracy.
5.1 Experimental Details
@@
Poisoning attacks. We provide the key implementation details below:
  1. $\bullet$
@@
  2. $\bullet$
Further implementation details are deferred to Appendix A.1.
Evaluating unlearning. When evaluating an unlearning method, a common hyperparameter across all the models is the compute budget (typically the number of gradient steps) given to the model. Clearly, if the compute budget is greater than that required for retraining the model from scratch, then the method is useless; Thus, the smaller the budget for a given level of performance the better. To put all the methods on equal footing, we allow each of them to use up to $10\%$ of the compute used in initial training (or fine-tuning) of the model (we also experiment with 4%, 6%, and 8% for comparison). This is actually quite generous, given that in modern settings like training a large language or vision model, $10\%$ of training compute is still significant in terms of time and cost; practical unlearning algorithms should ideally work with far less compute.
Computation aside, when evaluating the efficacy of an unlearning method two additional objectives are essential: validity of the unlearning process in that the algorithm effectively removes the forget set from the trained model, and model performance post-unlearning. For example, an unlearning algorithm that simply outputs a constant model might have removed the influence of the forget set, but it would not be very useful. We measure post-unlearning performance by comparing the test classification accuracy of the updated model to the model retrained without the poisoned data.
To gauge unlearning validity against different poisoning attacks, we use different metrics for targeted attacks, Gaussian poisons, and indiscriminate attacks.
  1. $\bullet$
    For indiscriminate data poisoning attacks, the goal is to decrease test accuracy, and so we can conclude that an unlearning algorithm is successful if the test accuracy after unlearning approaches that of a retrained model – note this is the same metric as for model performance.
  2. $\bullet$
    For targeted data poisoning attacks, where the goal is to cause the misclassification of a specific set of datapoints, an unlearning algorithm is valid if the misclassification rate on this specific set of datapoints is close to that of the retrained model. Note in this case that this is distinct from model performance, which measures test accuracy.
  3. $\bullet$
    For Gaussian data poisoning attacks, we first assess how good unlearning works by measuring how much information the Gaussian poisons leak from the model when no unlearning is performed, labeled as No unlearning in all figures. It represents the TPR at low FPR of the poisoned model before unlearning (solid orange lines in Figures 2 and 3). We then evaluate the success of the unlearning process by determining if the forget set is effectively removed and if the model’s original behavior is restored. Ideally, the TPR at low FPR should equal the FPR (dashed orange lines in Figure 2).
@@ 5.2 Experimental Results
Figure 2: Unlearning fails to remove Gaussian poisons across a variety of unlearning methods. We poison 1.5% of the training data by adding Gaussian noise with standard deviation $\varepsilon_{p,{\text{IMDb}}}^{2}=0.1$ and $\varepsilon_{p,\text{CIFAR-10}}^{2}=0.32$, respectively. We train/finetune a Resnet18 for 100 epochs and a GPT-2 for 10 epochs on the poisoned training datasets, respectively. Finally, we use $10\%$ of the original compute budget (i.e., 1 or 10 epochs) to unlearn the poisoned points. None of the unlearning methods removes the poisoned points as the orange vertical bars do not match the dashed orange retraining benchmark.
@@
Figure 3: Unlearning fails to remove targeted poisons across a variety of unlearning methods. We poison 1.5% of the training data by adding Witch’s Brew poisons (Geiping et al., 2021) to a Resnet-18 trained on CIFAR-10 or instruction poisons (Wan et al., 2023) to a GPT-2 finetuned on IMDb. We then train/finetune a Resnet-18 for 100 epochs and a GPT-2 for 10 epochs on the poisoned training datasets, respectively. In both cases, we use roughly $1/10$ of the original compute budget (10 epochs for CIFAR-10 or 1 epoch for IMDb) to unlearn the poisoned points. None of the considered methods remove the poisoned points.
Key observations and takeaways:
  1. $\bullet$
@@
    No silver bullet unlearning algorithm that can mitigate data poisoning. None of the evaluated methods completely remove the poisons from the trained models; See Figures 2, 3, and Table 2 and the caption therein for details on the failure of unlearning methods to remove poisons. The respective plots show that none of the methods performs on par with retraining from scratch in terms of post-unlearning test accuracy and effectiveness in removing the effects of data poisoning. Our experiments thus suggest that we need to develop better approximate unlearning methods for deep learning settings.
  2. $\bullet$
@@
    Different data poisoning methods introduce different challenges for unlearning. We observe that the success of an unlearning method in mitigating data poisoning depends on the poison type. For example, while GD can successfully alleviate the effects of indiscriminate data poisoning attacks for vision classification tasks, it typically fails to mitigate targeted or Gaussian poisoning attacks even while maintaining competitive model performance. Along similar lines, while SCRUB succeeds in somewhat mitigating Gaussian data poisoning in text classification tasks, it completely fails to mitigate targeted or indiscriminate data poisoning. This suggests that the different data poisoning methods complement each other and that to validate an unlearning algorithm, we need to consider all the above-mentioned data poisoning methods, along with other (preexisting) evaluations for unlearning.
  3. $\bullet$
@@
    The success of an unlearning method depends on the underlying task. We observe that various unlearning algorithms exhibit different behaviors for image classification and text classification tasks. For example, for data poisoning on a GPT-2 model, while EUk and NGD succeed in alleviating Gaussian data poisoning for the model trained with a classification head, they fail to remove targeted data poisoning on the same model trained with a text decoder. [^5] Our hypothesis for why EUk fails for text generation tasks is that it results in severe degradation of the model’s text generation capabilities due to re-initialization and fine-tuning of the last $k$ layers of the model from scratch. Similarly, GA succeeds in removing the effects of Gaussian and targeted data poisoning for Resnet-18 but fails to have a similar improvement for GPT-2 model. This suggests that the success of an approximate unlearning method over one task may not transfer to other tasks, and thus further research is needed to make transferable approximate unlearning approaches for deep learning settings.
Detailed comparison of different unlearning algorithms. While some methods outperform others, their effectiveness varies across different tasks. We mention our key observations below:
  1. $\bullet$
  2. $\bullet$
  3. $\bullet$
5.3 Ablation Studies
@@
Our ablation experiments, detailed in Appendix B, explore 1) the impact of varying the number of update steps and 2) the effect of varying the forgetset size. For methods like NGD, increasing the number of update steps generally enhances unlearning effectiveness (see Figure 8(b), orange bars). However, applying NGD to LLM models results in a substantial decrease in post-unlearning test accuracy, dropping by 10%. Conversely, for methods like EUk, additional steps do not improve unlearning or post-unlearning test accuracy (see Figure 8(a)). These trends are summarized in Figure 8. Furthermore, we experimented with different sizes of the forgetset. For Gaussian poisoning attacks, the results, summarized in Figures 10 and 9 of Appendix B, confirm consistent trends when 1.5%, 2%, and 2.5% of the training dataset are poisoned.
#Epochs Retrain NGP/GA GD CFk EUk SCRUB
@@ 6 Understanding why unlearning fails to remove poisons?
In Section 5.2, we demonstrated that various state-of-the-art approximate machine unlearning algorithms fail to fully remove the effects of data poisoning. Given these results, one may wonder what is special about the added poison samples, and why gradient-based unlearning algorithms fail to rectify their effects. In the following, we provide two hypotheses for understanding the failure of unlearning methods. We validate these hypotheses using a set of simple experiments based on linear and logistic regression on Resent-18 features which allow us to study these hypotheses experimentally. Thanks to the convexity of the corresponding loss the objectives have unique global minimizers making it easier to understand model shifts due to unlearning.
@@
Figure 4: Model shift for logistic regression on Resnet-18 features for CIFAR-10 dataset. The blue and the red curves denote the distance $\|\theta(S_{\textsf{corr}})-\theta(S_{\textsf{corr}}\setminus S_{\textsf{poison}}^{(\beta)})\|_{1}$, for indiscriminate and targeted data poisoning respectively, where $\beta$ denotes the corresponding percentage of poison samples that are unlearnt and for a dataset $S^{\prime}$, $\theta(S^{\prime})$ denotes a model trained from scratch on $S^{\prime}$. The orange curve plots the distance $\|\theta(S)-\theta(S\setminus S_{\mathrm{rand}}^{(\beta)})\|_{1}$ corresponding to randomly unlearning random clean training samples.
To validate this hypothesis, Figure 4 the $\ell_{1}$ shows the norm of the model shift introduced by unlearning (different amounts of) data poisons and random clean training data for logistic regression over feature representations given by the last layer of a fixed Resent-18 network (which corresponds to only updating the last layer of Resnet-18 model). Figure 4 shows that data poisons introduce much larger model shifts in the $\ell_{2}$ norm as compared to random training samples. We defer the experiment details to Appendix C.
Hypothesis 2: Poison samples shift the model in a subspace orthogonal to clean training samples. We next hypothesize that training with poison samples not only shifts the model by a larger amount, but the resultant shift lies in a subspace orthogonal to the span of clean training samples. Thus, gradient-based update algorithms that attempt unlearning with clean samples fail to counteract shifts within this orthogonal subspace and are unable to mitigate the impacts of data poisoning. We argue that to completely unlearn the effects of poison samples, an unlearning algorithm must incorporate gradient updates that specifically utilize these poison samples. However, employing a method like gradient ascent with poison samples is not ideal as it can degrade the overall performance of the model.
To validate this hypothesis, in Figure 5, we plot the inner product between the gradient update direction for gradient descent using clean training samples and the desired model shift direction, for unlearning for data poisons and random clean training samples respectively, for a simple linear regression task. The random subset of clean training samples is chosen so as to equate the model shift in both unlearning data poisons and random training samples. Figure 5 shows that the desired unlearning direction for data poisons is orthogonal to the update direction from gradient descent (as the respective cosine similarity between the two update directions is small). Experiment details are deferred to Appendix C.
@@ 7 Conclusion
Our experimental evaluation of state-of-the-art machine unlearning methods across different models and data modalities reveals significant shortcomings in their ability to effectively remove poisoned data points from a trained model. Despite various approaches which attempt to mitigate the effects of data poisoning, none were able to consistently approach the benchmark results of retraining the models from scratch. This highlights a critical gap in the true efficacy and thus practical value of current unlearning algorithms, questioning their validity in real-world applications where these unlearning methods may be deployed to ensure privacy, data integrity, or to correct model biases.
@@
Furthermore, our experiments demonstrate that the performance of unlearning methods varies significantly across different types of data poisoning attacks and models, indicating a lack of a one-size-fits-all solution. Given the increasing reliance on machine learning in critical and privacy-sensitive domains, our findings emphasize the importance of advancing rigorous research in machine unlearning to develop more effective, efficient, and trustworthy methods, that are either properly evaluated or have provable guarantees for unlearning. Future work should focus on creating novel unlearning algorithms that can achieve the dual goals of maintaining model integrity and protecting user privacy without the prohibitive costs associated with full model retraining.
Acknowledgements
AS acknowledges support from the Simons Foundation and NSF through award DMS-2031883, as well as from the DOE through award DE-SC0022199. GK is supported by a Canada CIFAR AI Chair, an NSERC Discovery Grant, and an unrestricted gift from Google.
Appendix A Implementation Details
A.1 Data Poisoning Attacks
A.1.1 Targeted Data Poisoning for Image Classification
$\min_{\Delta\in\Gamma}\ell\left(f(x_{\textsf{target}},\theta(\Delta)),y_{\mathrm{adv}}\right)\quad\text{where}$
where the constraint set $\Gamma\vcentcolon={}\left\{\Delta\mid{}\|\Delta(x)\|_{\infty}\leq\varepsilon_{p}\forall x\in S_{\textsf{poison}}\right\}.$ However, since directly solving (1) is computationally intractable due to its bi-level nature, Geiping et al. (2021) has opted for the approach to implicitly minimize the adversarial loss such that for any model $\theta$,
$\small\nabla_{\theta}(\ell(f(x_{\textsf{target}},\theta),y_{\textsf{advs}}))\approx\frac{\sum_{i=1}^{P}\nabla_{\theta}\ell\left(f(x^{i}+\Delta^{i},\theta),y^{i}\right)}{P}..$ (2)
1: Clean network $f(\cdot;\theta_{\mathrm{clean}})$ trained on uncorrupted base images $S_{\textsf{clean}}$ The target $(x_{\textsf{target}},y_{\textsf{target}})$ and the adversarial label $y_{\textsf{advs}}$ Poison budget $P$ and perturbation bound $\varepsilon_{p}$ Number of restarts $R$ and optimization steps $M$
3:for $r=1,\dots R$   restarts do
4:     Randomly initialize perturbations $\Delta$ s.t. $\|\Delta\|_{\infty}\leq\varepsilon_{p}$.
5:     for $k=1,\dots,M$   optimization steps do
  • Clean dataset $S_{\textsf{clean}}$ is the CIFAR-10 training set;
  • Set a Perturbation bound $\varepsilon_{p}$ of 16.
  • Generate $\Delta$ using the algorithm outlined in Algorithm 1
A.1.2 Targeted Data Poisoning for Language Sentiment Analysis
  • Clean dataset $S_{\textsf{clean}}$ is the IMDb reviews training set;
  • Set the maximum sequence length of the tokenizer to 128.
A.1.3 Indiscriminate Data Poisoning
For a given poison budget $b_{p}$ and perturbation bound $\varepsilon_{p}$, we generate the poison samples by following the Gradient Canceling (GC) procedure of Lu et al. (2023, 2024), a state-of-the-art indiscriminate poisoning attack in machine learning. In Gradient Canceling (GC) procedure, the adversary first finds a bad model $\theta_{\mathrm{low}}$ that has low-performance accuracy on the test dataset and then computes the perturbations $\Delta$ by solving the minimization problem
$\operatorname*{argmin}_{\Delta\in\Gamma}\leavevmode\nobreak\ \tfrac{1}{2}\|\widehat{\operatorname{\mathbb{E}}}_{(x,y)\in S_{\textsf{clean}}}\left[\nabla_{\theta}\ell((x,y);\theta_{\mathrm{low}})\right]+\widehat{\operatorname{\mathbb{E}}}_{(x,y)\in S_{\textsf{poison}}}\left[\nabla_{\theta}\ell((x+\Delta(x),y_{i});\theta_{\mathrm{low}}\right])\|_{2}^{2},$ (4)
where the constraint set $\Gamma\vcentcolon={}\left\{\Delta\mid{}\|\Delta(x)\|_{\infty}\leq\varepsilon_{p}\forall x\in S_{\textsf{poison}}\right\}.$ Informally speaking, the above objective function enforces that the generated poison points are such that $\theta_{\mathrm{low}}$ has vanishing (sub)gradients over the corrupted training dataset, and is thus close to a local minimizer of the training objective using the corrupted dataset. The model $\theta_{\mathrm{low}}$ is generated by the GradPC procedure of Sun et al. (2020), which is a gradient-based approach to finding a set of corrupted parameters that returns the lowest test accuracy within a certain distance from an input trained parameter. We provide the pseudocode of this attack in Algorithm 2.
Algorithm 2 Gradient Canceling (GC) Attack (Lu et al., 2023)
1: An uncorrupted clean dataset $S_{\textsf{clean}}$ Target network $f(\cdot;\theta_{\mathrm{low}})$ generated by GradPC (Sun et al., 2020) Poisoning budget $b_{p}$ and perturbation bound $\varepsilon_{p}$ Step size $\eta$
4:for $t=1,2,\dots$ do
6:     Calculate loss $\mathcal{L}=\leavevmode\nobreak\ \tfrac{1}{2}\|{g}_{c}+{g}_{p}\|_{2}^{2}$.
  • Clean dataset $S_{\textsf{clean}}$ is the CIFAR-10 training set;
  • Perturbation bound $\varepsilon_{p}$ is set to be infinite. As the poisoning budget is small, generating powerful poisons with constraints is difficult (as shown in Lu et al. (2023)). Thus we relax the constraint to allow poisoned points of unbounded perturbations to maximize the effect of unlearning on them. Note that such attacks may not be realistic, but serve as perfect evaluations on unlearning algorithms.
  • Target parameters $\theta_{\mathrm{low}}$ are generated by GradPC with a budget of $\varepsilon_{w}=1$, where $\varepsilon_{w}$ measures the L2 distance between $\theta_{\mathrm{low}}$ and the clean parameter.
A.1.4 Gaussian Data Poisoning

### v2

@@ 5 Can Machine Unlearning Remove Poisons?
We now evaluate state-of-the-art unlearning attacks for the task of removing both targeted and untargeted data poisoning attacks across vision and language models.
@@
(a) Standard MIA (Shokri et al., 2017)
(b) MIA via Gaussian poisoning
Figure 2: Standard MIA evaluations are insufficient for detecting unlearning violations. Left: At a low false positive rate (FPR=0.01), standard MIAs have low true positive rates, making them ineffective at identifying whether a targeted sample was successfully unlearned. Right: Our proposed Gaussian poison attack achieves a higher true positive rate at the same FPR, improving the detection of unlearning failures. A full trade-off curve comparison is provided in Figure 10.
[…]
Evaluating unlearning. When evaluating the efficacy of an unlearning method two objectives are essential: i) We measure post-unlearning performance by comparing the test classification accuracy of the updated model to the model retrained without the poisoned data. ii) To gauge unlearning validity against different poisoning attacks, we use different metrics for targeted attacks, Gaussian poisons, and indiscriminate attacks. For indiscriminate data poisoning attacks, the goal is to decrease test accuracy, and so we can conclude that an unlearning algorithm is successful if the test accuracy after unlearning approaches that of a retrained model – note this is the same metric as for model performance. For targeted data poisoning attacks, where the goal is to cause the misclassification of a specific set of datapoints, an unlearning algorithm is valid if the misclassification rate on this specific set of datapoints is close to that of the retrained model. Note in this case that this is distinct from model performance, which measures test accuracy. For Gaussian data poisoning attacks, we first assess how good unlearning works by measuring how much information the Gaussian poisons leak from the model when no unlearning is performed, labeled as No unlearning in all figures. It represents the TPR at low FPR of the poisoned model before unlearning. We then evaluate the success of the unlearning process by determining if the forget set is effectively removed and if the model’s original behavior is restored, labeled as Perfect Unlearning in all Figures.
5.1 Standard MIA Unlearning Evaluations can be Misleading
Prior work typically evaluates the efficacy of unlearning methods using MIAs (Shokri et al., 2017). However, Figure 2 shows that this approach is insufficient. None of the machine unlearning algorithms fully eliminate the influence of the deletion set $U$ from the updated model $\theta_{\textsf{updated}}$ when evaluated using our proposed Gaussian poisoning attack. Although most algorithms perform well against the standard MIA, this can lead to a misleading conclusion. An auditor relying solely on standard MIAs might incorrectly assume that all these methods effectively unlearn data (see Figure 2(a)). As demonstrated in Figure 2(b), Gaussian data poisoning reveals that approximate unlearning methods do not remove the points from set $U$, despite their success under the standard MIA.
@@ 5.2 Experimental Results
Figure 3: Unlearning fails to remove Gaussian poisons across a variety of unlearning methods. We poison 1.5% of the training data by adding Gaussian noise with standard deviation $\varepsilon_{p,{\text{IMDb}}}^{2}=0.1$ and $\varepsilon_{p,\text{CIFAR-10}}^{2}=0.32$, respectively. We train/finetune a Resnet18 for 100 epochs and a GPT-2 for 10 epochs on the poisoned training datasets, respectively. Finally, we use $10\%$ of the original compute budget (i.e., 1 or 10 epochs) to unlearn the poisoned points. None of the unlearning methods removes the poisoned points as the orange vertical bars do not match the dashed black retraining benchmark.
@@
Figure 4: Unlearning fails to remove targeted and backdoor poisons across a variety of unlearning methods. We poison 1.5% of the training data by adding Witch’s Brew poisons (Geiping et al., 2021) to a Resnet-18 trained on CIFAR-10 or instruction poisons (Wan et al., 2023) to a GPT-2 finetuned on IMDb. We then train/finetune a Resnet-18 for 100 epochs and a GPT-2 for 10 epochs on the poisoned training datasets, respectively. In both cases, we use roughly $1/10$ of the original compute budget (10 epochs for CIFAR-10 or 1 epoch for IMDb) to unlearn the poisoned points. None of the considered methods remove the poisoned points.
Below we discuss our key observations and main experimental takeaways and defer more detailed relative comparison between unlearning methods to Appendix E.1:
@@
1) No silver bullet unlearning algorithm that can mitigate data poisoning. None of the evaluated methods completely remove the poisons from the trained models; See Figures 3, 4, and Table 1 and the caption therein for details on the failure of unlearning methods to remove poisons. The respective plots show that none of the methods performs on par with retraining from scratch in terms of post-unlearning test accuracy and effectiveness in removing the effects of data poisoning, thus suggesting that we need to develop better approximate unlearning methods for deep learning settings.
@@
2) Different data poisoning methods introduce different challenges for unlearning. We observe that the success of an unlearning method in mitigating data poisoning depends on the poison type. For example, while GD can successfully alleviate the effects of indiscriminate data poisoning attacks for vision classification tasks, it typically fails to mitigate targeted or Gaussian poisoning attacks even while maintaining competitive model performance. Along similar lines, while SCRUB succeeds in somewhat mitigating Gaussian data poisoning in text classification tasks, it completely fails to mitigate targeted or indiscriminate data poisoning. This suggests that the different data poisoning methods complement each other and that to validate an unlearning algorithm, we need to consider all the above-mentioned data poisoning methods, along with other (preexisting) evaluations.
@@
3) The success of an unlearning method depends on the underlying task. We observe that various unlearning algorithms exhibit different behaviors for image classification and text classification tasks, e.g., for data poisoning on a GPT-2 model, while EUk and NGD succeed in alleviating Gaussian data poisoning for the model trained with a classification head, they fail to remove targeted data poisoning on the same model trained with a text decoder. [^3] Our hypothesis for why EUk fails for text generation is that it results in severe degradation of the model’s text generation capabilities due to re-initialization and fine-tuning of the last $k$ layers of the model from scratch. Similarly, GA succeeds in alleviating Gaussian and targeted data poisoning for Resnet-18 but fails to have a similar improvement for GPT-2 model. This suggests that the success of an approximate unlearning method over one task may not transfer to other tasks, and thus further research is needed to make transferable approximate unlearning methods.
5.3 Sensitivity Analysis and Additional Insights
@@
Additional experiments detailed in Appendix E, explore 1) the impact of varying the number of update steps and 2) the effect of varying the forgetset size. For methods like NGD, increasing the number of update steps generally enhances unlearning effectiveness (see Figure 11(b), orange bars). However, applying NGD to LLM models results in a substantial decrease in post-unlearning test accuracy, dropping by 10%. Conversely, for methods like EUk, additional steps do not improve unlearning or post-unlearning test accuracy (see Figure 11(a)). These trends are summarized in Figure 11. Furthermore, we experimented with different sizes of the forgetset. For Gaussian poisoning attacks, the results, summarized in Figures 13 and 12 of Appendix E, confirm consistent trends when 1.5%, 2%, and 2.5% of the training dataset are poisoned.
#Epochs Retrain NGD/NGP/GA GD CFk EUk SCRUB
@@ 6 Understanding why unlearning fails to remove poisons?
In Section 5.2, we demonstrated that various state-of-the-art approximate machine unlearning algorithms fail to fully remove the effects of data poisoning. Given these results, one may wonder what is special about the added poison samples, and why gradient-based unlearning algorithms fail to rectify their effects. In the following, we provide two hypotheses for understanding the failure of unlearning methods. We validate these hypotheses using a set of experiments based on linear and logistic regression on Resent-18 features which allow us to study these hypotheses experimentally. Thanks to the convexity of the corresponding loss the objectives have unique global minimizers making it easier to understand model shifts due to unlearning.
@@
Figure 5: Model shift for logistic regression on Resnet-18 features for CIFAR-10 dataset. The x-axis is the number of epochs. The blue and the red curves denote the distance $\|\theta(S_{\textsf{corr}})-\theta(S_{\textsf{corr}}\setminus S_{\textsf{poison}}^{(\beta)})\|_{1}$, for indiscriminate and targeted data poisoning respectively, where $\beta$ is the corresponding percentage of unlearned poison samples. For a dataset $S^{\prime}$, $\theta(S^{\prime})$ denotes a model trained from scratch on $S^{\prime}$. The orange curve plots the distance $\|\theta(S)-\theta(S\setminus S_{\mathrm{rand}}^{(\beta)})\|_{1}$ corresponding to unlearning random clean training samples.
To validate this hypothesis, Figure 5 shows the $\ell_{1}$ norm of the model shift introduced by unlearning data poisons and random clean training data, demonstrating that data poisons introduce much larger model shifts in this norm as compared to random training samples.
Hypothesis 2: Poison samples shift the model in a subspace orthogonal to clean training samples. We next hypothesize that training with poison samples not only shifts the model by a larger amount, but the resultant shift lies in a subspace orthogonal to the span of clean training samples. Thus, gradient-based update algorithms that attempt unlearning with clean samples fail to counteract shifts within this orthogonal subspace and are unable to mitigate the impacts of data poisoning. To completely unlearn the effects of poison samples, an unlearning algorithm must incorporate gradient updates that specifically utilize these poison samples. To validate this hypothesis, in Figure 14, we plot the inner product between the gradient update direction for gradient descent using clean training samples and the desired model shift direction, for unlearning for data poisons and random clean training samples respectively, for a simple linear regression task. The random subset of clean training samples is chosen so as to equate the model shift in both unlearning data poisons and random training samples. Figure 14 shows that the desired unlearning direction for data poisons is orthogonal to the update direction from gradient descent as the cosine similarity between the update directions is small.
@@ 7 Conclusion
Our experimental evaluation of state-of-the-art machine unlearning methods across different models and data modalities reveals significant shortcomings in their ability to effectively remove poisoned data points from a trained model. Despite various approaches which attempt to mitigate the effects of data poisoning, none were able to consistently approach the benchmark results of retraining the models from scratch. This highlights a critical gap in the true efficacy and thus practical value of current unlearning algorithms, questioning their validity in real-world applications where these unlearning methods may be deployed to ensure privacy, data integrity, or to correct model biases. Furthermore, our experiments demonstrate that the performance of unlearning methods varies significantly across different types of data poisoning attacks and models, indicating a lack of a one-size-fits-all solution. Given the increasing reliance on machine learning in critical and privacy-sensitive domains, our findings emphasize the importance of advancing rigorous research in machine unlearning to develop more effective, efficient, and trustworthy methods, that are either properly evaluated or have provable guarantees for unlearning. Future work should focus on creating novel unlearning algorithms that can achieve the dual goals of maintaining model integrity and protecting user privacy without the prohibitive costs associated with full model retraining.
Appendix A Additional Related Work
Reference Attack Types Model Architecture Unlearning Method
Marchant et al. (2022a) Backdoor attack Convex model Using Influence Functions (IF)
This work Indiscriminate Attack, Targeted Attack, Backdoor Attack, Gaussian Attack Any Any
Table 2: Comparing the data poisoning settings of this work to Marchant et al. (2022a).
Data poisoning attacks. In a data poisoning attack, an adversary may introduce or modify a small portion of the training data, and their goal is to elicit some undesirable behavior in a model trained on said data. One type of attack is a targeted data poisoning attack (Koh & Liang, 2017; Shafahi et al., 2018; Huang et al., 2020; Guo & Liu, 2020; Aghakhani et al., 2021), in which the goal is to cause a model to misclassify a specific point in the test set. Another type of attack is an untargeted (or indiscriminate) data poisoning attack (Biggio et al., 2012; Muñoz-González et al., 2017; Steinhardt et al., 2017; Koh et al., 2022; Lu et al., 2022; 2023), wherein the attacker seeks to reduce the test accuracy as much as possible. Though we do not focus on them in our work, there also exist backdoor attacks (Gu et al., 2017), in which training points are poisoned with a backdoor pattern, such that test points including the same pattern are misclassified and various detection techniques (Shan et al., 2022).
Poisoning machine unlearning systems. An orthogonal line of work investigates data poisoning attacks against machine unlearning pipelines (see, e.g., Chen et al. (2021); Marchant et al. (2022b); Carlini et al. (2022b); Di et al. (2023); Qian et al. (2023); Liu et al. (2024)). These works generally show that certain threats can arise even if unlearning is performed with provable guarantees, whereas we focus on data poisoning threats in standard (i.e., not machine unlearning) pipelines, that ought to be removed by an effective machine unlearning procedure (in particular, they would be removed by retraining from scratch without the poisoned points).
Evaluation works. Several prior works on machine unlearning evaluation have explored verifying the effect of unlearning through data poisoning in various settings and context (Wu et al., 2023; Marchant et al., 2022a; Sommer et al., 2020). In particular,  Wu et al. (2023) address the problem of graph unlearning, where they evaluate attacks involve adding adversarial edges to a graph. The authors demonstrate that both the influence function method and its extension, GIF, can mitigate the impact of these adversarial edges.
Meanwhile, Marchant et al. (2022a) focus exclusively on unlearning through Influence Functions (IF) on convex models. In this setting, the authors introduced a specifically designed backdoor data poisoning attack which is optimized knowing that the model owner will train, deploy, and updates their convex model using influence functions. Furthermore, the conclusion of their work - that the field of machine unlearning requires more rigorous evaluations - aligns with this work, considering the poisoning methods implemented in our work require less knowledge of the target model and are agnostic to the specific unlearning methods used (see Table 2).
Sommer et al. (2020) provide a framework for verifying exact data deletion. In a fundamentally different setting, the authors evaluate whether an MLaaS provider complies with a deletion request by completely removing the data point from the trained model. Hence, the findings in their work are independent of any unlearning methods. To verify a complete takedown, their framework involves running backdoor attacks to subsequently check if complete case-based deletion was done by the MLaaS provider.
Appendix B Gaussian Data Poisoning